# Interaction of the Primitive Equations with Sea Ice Dynamics

### v3

@@ 10. Global Well-Posedness close to Constant Equilibria
In this section, we verify that a slightly simplified version of the coupled system (3.4), or, equivalently, (4.6), is globally well-posed for initial data close to constant equilibria. To this end, we employ the so-called generalized principle of linearized stability, see [42] or [41, Section 5.3].
@@
Similarly as in Subsection 8.2, see also [9, Sections 6 and 7], we verify that
@@
The next step is to verify that an equilibrium $u_{*}$ of the above shape is normally stable in the sense of [41, Theorem 5.3.1]. To this end, we first investigate the spectral properties of $\mathcal{A}_{0}$.
[…]
@@
Next, from [9, (4.4) and (7.4)] we recall that
@@
(10.7) $-\langle{\Delta_{\mathrm{H}}h},{h}\rangle_{\mathrm{L}^{2}(G)}=\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}\quad\text{and}\quad-\langle{\Delta_{\mathrm{H}}a},{a}\rangle_{\mathrm{L}^{2}(G)}=\|\nabla_{\mathrm{H}}a\|_{\mathrm{L}^{2}(G)}^{2}.$
@@
(10.8) $({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},c_{3}{v}_{\mathrm{ice}},c_{4}h,c_{5}a),\quad\text{where}\quad c_{3}=\frac{1}{C_{\mathrm{o,i}}(h_{*})},\enspace c_{4}=c_{3}\frac{\partial_{h}P_{*}}{2{\rho}_{\mathrm{ice}}h_{*}^{2}}\text{ and }c_{5}=c_{3}\frac{\partial_{a}P_{*}}{2{\rho}_{\mathrm{ice}}h_{*}a_{*}}.$
@@
00 0 $=\lambda\|{v}_{\mathrm{atm}}\|_{{\mathrm{L}}_{\overline{\sigma}}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\lambda\|{v}_{\mathrm{ocn}}\|_{{\mathrm{L}}_{\overline{\sigma}}^{2}({\Omega}_{\mathrm{ocn}})}^{2}+\lambda c_{3}\|{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+\lambda c_{4}\|h\|_{\mathrm{L}^{2}(G)}^{2}+\lambda c_{5}\|a\|_{\mathrm{L}^{2}(G)}^{2}$
@@
where $\vec{n}$ denotes the outer normal vector. For the oceanic term, it is crucial to note that $\vec{n}=(0,0,\pm 1)^{\top}$ on the upper and lower boundary, so the coupling condition with sea ice on $\Gamma_{o}$ does not come into play. For a similar result, we also refer to [26, Lemma 6.3]. By $({u}\cdot\nabla){v}=({v}\cdot\nabla_{\mathrm{H}}){v}+w({v})\cdot\partial_{z}{v}$, we then get
@@
We mainly write $\varepsilon=\varepsilon({v}_{\mathrm{ice}})$ in the sequel for convenience. For $u\in V$, we especially deduce that $P(h,a)\geq p^{*}\kappa\exp{(-c)}=:P_{**}$ as well as $\frac{1}{\triangle_{\delta}(\varepsilon)}\geq\frac{1}{\sqrt{\delta+c_{e}r^{2}}}$ for a suitable constant $c_{e}>0$, see also the proof of Lemma 7.2 in [9]. We emphasize that $P_{**}$ is independent of $u$, $\delta$ and $r$. Using an integration by parts in conjunction with the periodic boundary conditions, exploiting the previous estimates, invoking $\varepsilon^{\top}\mathbb{S}\varepsilon=\triangle^{2}(\varepsilon)\geq 0$ by virtue of [9, (4.4)] and employing Korn’s inequality, we infer that
@@
$-\langle{\partial_{\nu_{\Gamma_{o}}}{v}_{\mathrm{ocn}}},{{v}_{\mathrm{ice}}}\rangle_{\mathrm{L}^{2}(G)}+c_{3}{\mu}_{\mathrm{ocn}}\langle{\partial_{\nu_{\Gamma_{o}}}{v}_{\mathrm{ocn}}},{{v}_{\mathrm{ice}}}\rangle_{\mathrm{L}^{2}(G)}=0$
@@
(10.12) 00 0 $\geq\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+C_{1}\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}+\frac{c_{3}C_{K}}{\sqrt{\delta+c_{e}r^{2}}}\|\nabla_{\mathrm{H}}{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+c_{4}d_{\mathrm{h}}\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}$
@@
$\Big{\|}c_{3}\frac{\partial_{a}P(h,a)}{2}-c_{5}a\Big{\|}_{\mathrm{L}^{\infty}(G)}$ $=\frac{c_{3}}{2}\Big{\|}cp^{*}h\exp(-c(1-a))-cp^{*}h_{*}\frac{a}{a_{*}}\exp(-c(1-a_{*}))\Big{\|}_{\mathrm{L}^{\infty}(G)}$ $\leq C_{*}r.$
@@
00 0 $\geq\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\left(C_{1}-2C_{**}r-C_{***}(r+h_{*})r\right)\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}$
@@
00 0 $\geq C\left(\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}+\|\nabla_{\mathrm{H}}{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}+\|\nabla_{\mathrm{H}}a\|_{\mathrm{L}^{2}(G)}^{2}\right)$
@@
[…]
For the total linearization $\mathcal{A}_{0}$ with domain $D(\mathcal{A}_{0})=\mathcal{X}_{1}$, it holds that 00 is a semi-simple eigenvalue of $\mathcal{A}_{0}$, meaning that $N(\mathcal{A}_{0})\oplus R(\mathcal{A}_{0})=X_{0}$.
@@
First, the structure of $\mathcal{A}_{0}$ from (10.3) yields that $\mathrm{A}^{\mathrm{atm}}$ can be investigated separately. Using the representation of the hydrostatic Stokes operator with Neumann boundary conditions on the upper and lower boundary $\mathrm{A}^{\mathrm{atm}}$ from the proof of Theorem 3.1 in [16], we find that $\mathrm{A}^{\mathrm{atm}}$ inherits the property that 00 is a semi-simple eigenvalue from the respective Neumann-Laplacian operator. It is thus sufficient to consider the remaining matrix $\mathcal{A}_{0}^{\mathrm{rem}}$ given by
@@
when testing the equation $\mathcal{A}_{0}^{\mathrm{rem}}({v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)=0$ for $({v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)\in D(\mathcal{A}_{0,r}^{\mathrm{rem}})=D(\mathcal{A}_{0}^{\mathrm{rem}})\cap\mathcal{X}_{0,r}^{\mathrm{rem}}$ suitably. We deduce therefrom that ${v}_{\mathrm{ocn}}=0$ and ${v}_{\mathrm{ice}}$, $h$ as well as $a$ are constant so that $\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}={v}_{\mathrm{ice}}$ on $\Gamma_{o}$ and $h$, $a\in\mathrm{L}^{q}_{0}(G)$ yield that ${v}_{\mathrm{ice}}=0$ as well as $h=a=0$. Consequently, 00 is not an eigenvalue of $\mathcal{A}_{0,r}^{\mathrm{rem}}$. The compact resolvent, see Subsection 8.1(g), which carries over to the present setting, then implies that $0\in\rho(\mathcal{A}_{0,r}^{\mathrm{rem}})$.
@@
[…]
For $u_{*}$ as above, $-\mathcal{A}(u_{*})$ has the property of maximal regularity on $\mathcal{X}_{0}$.
@@
Combining the triangular structure of $\mathcal{A}(u_{*})$ as in (10.2) and the maximal regularity of $\mathrm{A}^{\mathrm{atm}}$, see [16, Corollary 3.4], and of the horizontal Laplacians on $G$, we find that the task reduces to establishing maximal regularity of
@@
As a result of the above argument in conjunction with Subsection 8.1(c), there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, it holds that $-\mathcal{A}_{\mathrm{o,i}}(u_{*})+\omega$ has the property of maximal regularity. It can be shown that $\omega_{0}$ can be chosen to be equal to the spectral bound of $s(\mathcal{A}_{\mathrm{o,i}}(u_{*}))$ of $\mathcal{A}_{\mathrm{o,i}}(u_{*})$. With regard to the $q$-independence of the spectrum, see Subsection 8.1, carrying over to the present situation, it suffices to study the spectrum in the $\mathrm{L}^{2}$-case.
@@
Concatenating Section 10, Section 10, Section 10 and Section 10, we find that an equilibrium $u_{*}$ of the above shape is normally stable. The generalized principle of linearized stability, see [41, Theorem 5.3.1] and also [42], yields the assertion of Theorem 3.2. ∎
Acknowledgements The first author would like to thank DFG for support through project 538212014. The second and third author gratefully acknowledge the support by DFG project FOR 5528.

### v4

@@ 10. Global Well-Posedness close to Constant Equilibria
In this section, we verify that a slightly simplified version of the coupled system (3.4), or, equivalently, (4.6), is globally well-posed for initial data close to constant equilibria. To this end, we employ the so-called generalized principle of linearized stability, see [35, Section 5.3].
@@
Similarly as in Subsection 8.2, see also [8, Sections 6 and 7], we verify that
@@
The next step is to verify that an equilibrium $u_{*}$ of the above shape is normally stable in the sense of [35, Theorem 5.3.1]. To this end, we first investigate the spectral properties of $\mathcal{A}_{0}$.
Lemma 10.1.
@@
Next, from [8, (4.4) and (7.4)] we recall that
@@
(10.7) $\langle{\Delta_{\mathrm{H}}h},{h}\rangle_{\mathrm{L}^{2}(G)}=\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}\quad\text{and}\quad-\langle{\Delta_{\mathrm{H}}a},{a}\rangle_{\mathrm{L}^{2}(G)}=\|\nabla_{\mathrm{H}}a\|_{\mathrm{L}^{2}(G)}^{2}.$
@@
(10.8) $({v}_{\mathrm{atm}},{v}_{\mathrm{ocn}},c_{3}{v}_{\mathrm{ice}},c_{4}h,c_{5}a),\quad\text{where}\quad c_{3}=\frac{1}{C_{\mathrm{o,i}}(h_{*})},\kern 5.0ptc_{4}=c_{3}\frac{\partial_{h}P_{*}}{2{\rho}_{\mathrm{ice}}h_{*}^{2}}\text{ and }c_{5}=c_{3}\frac{\partial_{a}P_{*}}{2{\rho}_{\mathrm{ice}}h_{*}a_{*}}.$
@@
$0$ $=\lambda\|{v}_{\mathrm{atm}}\|_{{\mathrm{L}}_{\overline{\sigma}}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\lambda\|{v}_{\mathrm{ocn}}\|_{{\mathrm{L}}_{\overline{\sigma}}^{2}({\Omega}_{\mathrm{ocn}})}^{2}+\lambda c_{3}\|{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+\lambda c_{4}\|h\|_{\mathrm{L}^{2}(G)}^{2}+\lambda c_{5}\|a\|_{\mathrm{L}^{2}(G)}^{2}$
@@
where $\vec{n}$ denotes the outer normal vector. For the oceanic term, it is crucial to note that $\vec{n}=(0,0,\pm 1)^{\top}$ on the upper and lower boundary, so the coupling condition with sea ice on $\Gamma_{o}$ does not come into play. For a similar result, we also refer to [25, Lemma 6.3]. By $({u}\cdot\nabla){v}=({v}\cdot\nabla_{\mathrm{H}}){v}+w({v})\cdot\partial_{z}{v}$, we then get
@@
We mainly write $\varepsilon=\varepsilon({v}_{\mathrm{ice}})$ in the sequel for convenience. For $u\in V$, we especially deduce that $P(h,a)\geq p^{*}\kappa\exp{(-c)}=:P_{**}$ as well as $\frac{1}{\triangle_{\delta}(\varepsilon)}\geq\frac{1}{\sqrt{\delta+c_{e}r^{2}}}$ for a suitable constant $c_{e}>0$, see also the proof of Lemma 7.2 in [8]. We emphasize that $P_{**}$ is independent of $u$, $\delta$ and $r$. Using an integration by parts in conjunction with the periodic boundary conditions, exploiting the previous estimates, invoking $\varepsilon^{\top}\mathbb{S}\varepsilon=\triangle^{2}(\varepsilon)\geq 0$ by virtue of [8, (4.4)] and employing Korn’s inequality, we infer that
@@
$\langle{\partial_{\nu_{\Gamma_{o}}}{v}_{\mathrm{ocn}}},{{v}_{\mathrm{ice}}}\rangle_{\mathrm{L}^{2}(G)}+c_{3}{\mu}_{\mathrm{ocn}}\langle{\partial_{\nu_{\Gamma_{o}}}{v}_{\mathrm{ocn}}},{{v}_{\mathrm{ice}}}\rangle_{\mathrm{L}^{2}(G)}=0$
@@
(10.12) $0$ $\geq\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+C_{1}\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}+\frac{c_{3}C_{K}}{\sqrt{\delta+c_{e}r^{2}}}\|\nabla_{\mathrm{H}}{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+c_{4}d_{\mathrm{h}}\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}$
@@
$\Big\|c_{3}\frac{\partial_{a}P(h,a)}{2}-c_{5}a\Big\|_{\mathrm{L}^{\infty}(G)}$ $=\frac{c_{3}}{2}\Big\|cp^{*}h\exp(-c(1-a))-cp^{*}h_{*}\frac{a}{a_{*}}\exp(-c(1-a_{*}))\Big\|_{\mathrm{L}^{\infty}(G)}\hskip-10.00002pt$ $\leq C_{*}r.$
@@
$0$ $\geq\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\left(C_{1}-2C_{**}r-C_{***}(r+h_{*})r\right)\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}$
@@
$0$ $\geq C\left(\|\nabla{v}_{\mathrm{atm}}\|_{\mathrm{L}^{2}({\Omega}_{\mathrm{atm}})}^{2}+\|{v}_{\mathrm{ocn}}\|_{\mathrm{H}^{1}({\Omega}_{\mathrm{ocn}})}^{2}+\|\nabla_{\mathrm{H}}{v}_{\mathrm{ice}}\|_{\mathrm{L}^{2}(G)}^{2}+\|\nabla_{\mathrm{H}}h\|_{\mathrm{L}^{2}(G)}^{2}+\|\nabla_{\mathrm{H}}a\|_{\mathrm{L}^{2}(G)}^{2}\right)$
@@
Lemma 10.3.
For the total linearization $\mathcal{A}_{0}$ with domain $D(\mathcal{A}_{0})=\mathcal{X}_{1}$, it holds that 0 is a semi-simple eigenvalue of $\mathcal{A}_{0}$, meaning that $N(\mathcal{A}_{0})\oplus R(\mathcal{A}_{0})=X_{0}$.
@@
First, the structure of $\mathcal{A}_{0}$ from (10.3) yields that $\mathrm{A}^{\mathrm{atm}}$ can be investigated separately. Using the representation of the hydrostatic Stokes operator with Neumann boundary conditions on the upper and lower boundary $\mathrm{A}^{\mathrm{atm}}$ from the proof of Theorem 3.1 in [17], we find that $\mathrm{A}^{\mathrm{atm}}$ inherits the property that 0 is a semi-simple eigenvalue from the respective Neumann-Laplacian operator. It is thus sufficient to consider the remaining matrix $\mathcal{A}_{0}^{\mathrm{rem}}$ given by
@@
when testing the equation $\mathcal{A}_{0}^{\mathrm{rem}}({v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)=0$ for $({v}_{\mathrm{ocn}},{v}_{\mathrm{ice}},h,a)\in D(\mathcal{A}_{0,r}^{\mathrm{rem}})=D(\mathcal{A}_{0}^{\mathrm{rem}})\cap\mathcal{X}_{0,r}^{\mathrm{rem}}$ suitably. We deduce therefrom that ${v}_{\mathrm{ocn}}=0$ and ${v}_{\mathrm{ice}}$, $h$ as well as $a$ are constant so that $\mathrm{tr}_{\Gamma_{o}}{v}_{\mathrm{ocn}}={v}_{\mathrm{ice}}$ on $\Gamma_{o}$ and $h$, $a\in\mathrm{L}^{q}_{0}(G)$ yield that ${v}_{\mathrm{ice}}=0$ as well as $h=a=0$. Consequently, 0 is not an eigenvalue of $\mathcal{A}_{0,r}^{\mathrm{rem}}$. The compact resolvent, see Subsection 8.1(g), which carries over to the present setting, then implies that $0\in\rho(\mathcal{A}_{0,r}^{\mathrm{rem}})$.
@@
Lemma 10.4.
For $u_{*}$ as above, $\mathcal{A}(u_{*})$ has the property of maximal regularity on $\mathcal{X}_{0}$.
@@
Combining the triangular structure of $\mathcal{A}(u_{*})$ as in (10.2) and the maximal regularity of $\mathrm{A}^{\mathrm{atm}}$, see [17, Corollary 3.4], and of the horizontal Laplacians on $G$, we find that the task reduces to establishing maximal regularity of
@@
As a result of the above argument in conjunction with Subsection 8.1(c), there is $\omega_{0}\in\mathbb{R}$ such that for all $\omega>\omega_{0}$, it holds that $\mathcal{A}_{\mathrm{o,i}}(u_{*})+\omega$ has the property of maximal regularity. It can be shown that $\omega_{0}$ can be chosen to be equal to the spectral bound of $s(\mathcal{A}_{\mathrm{o,i}}(u_{*}))$ of $\mathcal{A}_{\mathrm{o,i}}(u_{*})$. With regard to the $q$-independence of the spectrum, see Subsection 8.1, carrying over to the present situation, it suffices to study the spectrum in the $\mathrm{L}^{2}$-case.
@@
Concatenating Section 10, Section 10, Section 10 and Section 10, we find that an equilibrium $u_{*}$ of the above shape is normally stable. The generalized principle of linearized stability, see [35, Theorem 5.3.1], yields the assertion of Theorem 3.4. ∎
Acknowledgements. Tim Binz would like to thank DFG for support through project 538212014. Felix Brandt gratefully acknowledges the support by the German National Academy of Sciences Leopoldina through the Leopoldina Fellowship Program with grant number LPDS 2024-07. Felix Brandt and Matthias Hieber would like to thank DFG for support through project FOR 5528.